OpenStudy (anonymous):

What is the limit as x goes to zero for: ((sin(3x))/((sin(6x))?

myininaya (myininaya):

3/6=1/2

OpenStudy (anonymous):

ok, that was straight fwd; I got it right!

OpenStudy (anonymous):

I think :-)

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\sin(3x)}{3x}*\frac{6x}{\sin(6x)}*\frac{3}{6}\]

OpenStudy (anonymous):

thanks, that makes sense!

myininaya (myininaya):

1*1*3/6

myininaya (myininaya):

:)

OpenStudy (anonymous):

Or you could do: \[\lim_{x \rightarrow 0}\frac{\sin(3x)}{2\sin(3x)\cos(3x)}=\lim_{x \rightarrow 0}\frac{1}{2\cos(3x)}=\frac{1}{2\cos(0)}=\frac{1}{2}\].

OpenStudy (anonymous):

Using the double angle formula: sin(2x)=2sin(x)cos(x)